For a normal random variable, with  μ = 20, and  σ = 2, find the following probabilities.
(a) Pr(X  ≤  21.1)
(b) Pr(X  >  15.90)
Answer #1

Given u = 20 o=2 The required probability is, P(X$211)=P(434-21.1-4) =P[221.1-20) = P(Z <0.55) =0.7088 (From the standard norb) The required probability is, P(X>15.90)=P(X-4, 15.90 – 4 =PZ, 15.90-20 = P(Z >-2.05) =1-P(ZS-2.05) =1-0.0202 = 0.9798 (Fro